Exercise 5 Let 2 1 -2 -1 (a) Find A4 (b) Find eA (c) Find a matrix B such that B2 = A.
Answer #1

Pegè No. Date A= 2 -2 2 1: 2 A 2-1 -2 2 ASAA A A=A=A AAA A A- AA AA 2 2. 4 t AA 4-1 3 A, KEN It AtA 31 2 ) A + 4T 2 e IteA A=